On a Class of Set-Valued Variational Inclusion in Banach Spaces

Minru Bai ^{∗}

Department of Mathematics, Hunan University, Changsha,Hunan410082,P.R.China

Abstract— The purpose of this paper is to study a class of set-valued variational inclusions in Banach spaces. An iterative algorithm for solving this kind of setvalued variational incusions is established. We prove the existence of solutions and the convergence of iterative sequences generated by the algorithms in Banach spaces.

∗The work is supported by NNSF of China (10071017)